OpenStudy (anonymous):

WiIL MEDAL AND FAN! A right rectangluar prism has a square base and a height of 10 feet. It's volume, V= 90 cunic feet. What is the length of the sides of the base? Show your work.

OpenStudy (anonymous):



OpenStudy (anonymous):

Calcutlate the volume of the complex solid. Show your work. 

OpenStudy (anonymous):

Let side of base=x area of base=x^2 volume=base area x height 90=(x^2)*10 x= \[\sqrt{9}\] x=3